# NewReff.bib

### v1

@@ 1 Introduction
Discrete measured groupoids were introduced by Mackey in [Mackey62] and provide a uniform framework to study discrete countable groups, countable measured equivalence relations and (measurable) actions of countable discrete groups on measure spaces. In [Hahn78], Hahn studied the left regular representation of such groupoids and introduced their von Neumann algebras. Measured groupoids are of interest from an operator algebraic point of view because they give a huge class of examples of von Neumann algebras, that covers group von Neumann algebras, von Neumann algebras with Cartan subalgebras and group measure space von Neumann algebras. Every discrete measured groupoid comes equipped with an associated countable measured equivalence relation and a measured field of isotropy groups.
As in the case of groups and group actions, several properties of the associated von Neumann algebra is completely captured by the groupoid. In [BCDK24] we introduced the infinite conjugacy class (icc) property for groupoids and proved that the von Neumann algebra associated to a discrete measured groupoid is a factor if and only if the groupoid is icc and satisfies an ergodicity condition. In some sense, the conditions as well as the proof of the main theorem in [BCDK24] resemble the cases of groups and group actions. Moreover, the literature on discrete measured groupoids lacks concrete examples beyond equivalence relations and transformation groupoids. This naturally leads to asking how much more general the class of ergodic groupoids is than the sub-classes of ergodic equivalence relations and transformation groupoids of ergodic group actions. More precisely, are there ergodic discrete measured groupoids that are not isomorphic to equivalence relations or transformation groupoids?
This leads to the natural definition of genuine discrete measured groupoids. We call an ergodic groupoid genuine if it is not isomorphic to a transformation groupoid or an equivalence relation. More generally a discrete measured groupoid is called genuine if its direct integral decomposition into ergodic groupoids (as in [Hahn78]) has a positive measure part consisting of ergodic genuine groupoids. This article deals with the following question.
@@
The main idea comes from the following easy observation: If $G\ltimes X$ is the transformation groupoid of a non-singular group action $G\curvearrowright(X,\mu)$, the isotropy group at a point $x\in X$ is the stabilizer $\{g\in G\;|\;gx=x\}$ of the action, and in particular a subgroup of $G$. Thus if we can construct a measured field of groups $(G_{x})_{x\in X}$ such that they cannot be subgroups of a single discrete countable group, then such a measured field of groups cannot be the isotropy part of a transformation groupoid. To construct such a measured field, we use B.H. Neumann’s construction in [Neumann37] of an uncountable family of 2-generated groups that are pairwise non-isomorphic.
The interesting point about Neumann’s construction is that he associates very canonically to every strictly increasing sequence of odd integers $U$, a countable set of elements $X_{U}$ and two permutations $\alpha_{U}$ and $\beta_{U}$ of this countable set which makes the family of groups $G_{U}$ generated by these permutations mutually non-isomorphic. We can use the Borel bijection between $(0,1)$ and $\mathbb{N}^{\mathbb{N}}$ to get a standard Borel structure on the set $Y$ of all strictly increasing sequences of odd integers. Due to the canonical nature of the construction, we can endow $(G_{U})_{U\in Y}$ with a standard Borel structure that makes it a Borel field of groups as in Corollary 3.4. Now, by equipping $Y$ with any diffuse Borel probability measure, we get a measured field of groups which can not be the isotropy part of any transformation groupoid.
@@ 2.1 Discrete measured groupoids
The most basic example of a discrete measured groupoid is a countable discrete group with a unit space consisting of a single point. More generally for any such groupoid and every $x\in X$, the isotropy group at $x$ is the discrete countable group $\Gamma_{x}=\{g\in\mathcal{G}\;|\;s(g)=t(g)=x\}$. The subgroupoid $\operatorname{Iso}(\mathcal{G})$ given by $\{g\in\mathcal{G}\>|\;s(g)=t(g)\}$ is called the isotropy subgroupoid of $\mathcal{G}$. The isotropy subgroupoid consists of a measured family of discrete countable groups $(\Gamma_{x})_{x\in X}$, which can be made precise as follows:
@@
    If $\eta$ is a Borel probability measure on $Z$, then $(G_{z})_{z\in Z}$ is called a measured field of discrete countable groups over $(Z,\eta)$ if there is a $\eta$-conull subset $Z_{0}$ such that with the restricted Borel structure on $\pi^{-1}(Z_{0})$, $(G_{z})_{z\in Z_{0}}$ is a Borel field of discrete countable groups.
We remark here that while defining a Borel fields of Polish groups $K=(K_{x})_{x\in X}$ in the traditional way (see [Sutherland85]), one also requires the existence of enough Borel sections, i.e., a sequence of Borel sections $\theta_{n}:X\rightarrow K$ such that at each $x\in X$, $\{\theta_{n}({x})\;|\;n\in\mathbb{N}\}$ is dense in $K_{x}$. Such a sequence of sections is called a total set of sections. However note that in the case of discrete countable groups, this condition is automatic, essentially due to the Lusin-Novikov uniformization theorem (see [Kec95, Theorem 18.10]).
@@
Let $(Z,\eta)$ be a standard probability space and let $(G_{z})_{z\in Z}$ be a measured field of discrete countable groups. Let $K$ be a discrete countable group. Then the wreath products $(G_{x}\wr K)_{x\in X}$ form a measured field of discrete countable groups.
@@
In this case the projections $\pi_{s}$ and $\pi_{t}$ sending an element $(x,y)$ to $y$ and $x$ respectively are countable-to-one maps and send Borel sets to Borel sets. It turns out that the non-singularity condition above is equivalent to $\mu_{s}$ and $\mu_{t}$ being equivalent measures on $\mathcal{R}$. Any discrete measured groupoid $\mathcal{G}$ on $(X,\mu)$ comes equipped with such a countable measured equivalence relation $\mathcal{R}_{\mathcal{G}}$ given by:
@@
Orbit equivalence relations are the most basic examples of equivalence relations. Let $G$ be a countable discrete group and $G\curvearrowright(X,\mu)$ be a nonsingular action of $G$ on a standard probability space. Then the equivalence relation $\mathcal{R}_{G\curvearrowright X}$ given by $\{(gx,x)\;|\;g\in G,x\in X\}$ is a countable measured equivalence relation called the orbit equivalence relation. As shown in [Feldman-Moore-1, Theorem 1], every countable measured equivalence relation is in fact isomorphic to an orbit equivalence relation of a countable group action.
With any such action, one can also associate its transformation groupoid which intersects in some situations with the orbit equivalence relations as we see now. Once again let $G\curvearrowright(X,\mu)$ be a non-singular action on a standard probability space and consider the groupoid $X\rtimes G=\{(x,g)\;|\;x\in X,g\in G\}$ with unit space $(X,\mu)$ such that $s(x,g)=x$ and $t(x,g)=gx$. Multiplication and inverses are defined naturally as follows for all $g,h\in G$ and $x\in X$:
$(gx,h)(x,g)=(x,hg)\text{ and }(g,x)^{-1}=(gx,g^{-1})$
The isotropy group $\Gamma_{x}$ at $x\in X$ is then isomorphic to the stabilizer $\{g\in G\;|\;gx=x\}$. If the stabilizer groups are trivial almost everywhere, i.e. when the action is essentially free, then $X\rtimes G$ is isomorphic to the orbit equivalence relation as a discrete measured groupoid, where the isomorphism is given by $(x,g)\mapsto(gx,x)$. However the converse is not true, i.e., every countable measured equivalence relation does not arise as the orbit equivalence relation of an essentially free action, and is hence not isomorphic to a transformation groupoid, as shown by Furman in [Furman99].
@@ 2.2 Equivalence relations and ergodic decomposition
A lot of facts in this subsection are taken from [Feldman-Moore-1] and we shall state them without proofs here. Suppose that $\mathcal{R}$ is a countable measured equivalence relation on a standard probability space $(X,\mu)$. Recall that such a countable measured equivalence relation is called ergodic if for every Borel subset $E\subset X$ satisfying $\mathcal{R}(E)=E$ up to measure zero, $E$ is either null or co-null. Recall that the von Neumann algebra $L^{\infty}(X,\mu)^{\mathcal{R}}$ is the space of all $\mathcal{R}$-invariant bounded measurable functions on $(X,\mu)$ where two such functions are identified if they are equal almost everywhere. It can be checked that $\mathcal{R}$ is ergodic if and only if it is equal to $\mathbb{C}\cdot 1$.
Countable measured equivalence relations are broadly classified into three types, analoguous to the classification of von Neumann algebras. Let $\mathcal{R}$ be an ergodic equivalence relation on a standard probability space $(X,\mu)$. Then $\mathcal{R}$ is said to be of type In for $n\in\{1,2,...,\infty\}$ if $\mathcal{R}$ is isomorphic to the complete equivalence relation on a set of $n$ elements. If $\mathcal{R}$ is not necessarily ergodic, then $\mathcal{R}$ is said to be of type In if $\mathcal{R}$ is isomorphic to an equivalence relation on $E\times S$, where $E$ is a standard Borel space and $S$ is a set of $n$ elements given by $(x,k)\sim(y,l)$ if and only $x=y$. An equivalence relation is said to be smooth or discrete or simply of type I if $X$ admits a $\mathcal{R}$-invariant Borel partition $X=\bigsqcup_{n}X_{n}$ such that $\mathcal{R}|_{X_{n}}$ is of type In for $n\in\{1,2,...\infty\}$.
[…]
@@
Type I equivalence relations are precisely the ones which admit a so-called fundamental domain. More precisely we have the following characterization. One can look at standard texts or lecture notes on Borel equivalence relations for a proof, for example see [CalderoniLectureNotes09, Chapter 2]
@@
By the ergodic decomposition theorem ([Hahn78, Theorem 6.1]), every discrete measured groupoid is isomorphic to a direct integral of ergodic discrete measured groupoids. More precisely, given a discrete measured groupoid $\mathcal{G}$, let $\mathcal{G}=\mathcal{G}_{\text{s}}\bigsqcup\mathcal{G}_{\text{ns}}$ be its decomposition into its smooth and non-smooth parts. Then $\mathcal{G}_{\text{ns}}$ is isomorphic to a direct integral of ergodic non-smooth groupoids and $\mathcal{G}_{\text{s}}$ admits a decomposition $\mathcal{G}_{\text{s}}=\bigsqcup_{n\in\mathbb{N}}\mathcal{G}_{\text{s}}^{n}$, where $\mathcal{G}_{\text{s}}^{n}$ has associated equivalence relation of type In and $\mathcal{G}_{\text{s}}^{n}$ is a direct integral of ergodic discrete measured groupoids of type In. The ergodic decomposition of discrete measured groupoids is furthermore unique in some sense. We don’t go into the details here and refer the reader to [Hahn78]. This motivates the following main definition of this article.
@@ 2.3 Semi-direct product groupoids
Suppose that we have a countable measured equivalence relation $\mathcal{R}$ on a standard probability space $(X,\mu)$ and suppose $G=(G_{x})_{x\in X}$ is a measured field of discrete countable groups. Then an action of $\mathcal{R}$ on $G$ is given by a family of group isomorphisms $\delta_{(y,x)}:\Gamma_{x}\rightarrow\Gamma_{y}$ for all $(y,x)\in\mathcal{R}$ such that outside a null set we have:
@@
$(h,(z,y))\circ(g,(y,x))$ $=(\delta_{(x,y)}(z)\cdot g,(z,x))$
@@
It was shown in [Popa-Shlyakhtenko-Vaes20, Proposition 6.5] that for every so-called treeable equivalence relation $\mathcal{R}$, any discrete measured groupoid with associated equivalence relation $\mathcal{R}$ can be written as such a semi-direct product. In particular, this is the case if $\mathcal{R}$ is a hyperfinite equivalence relation, i.e., one that can be written as a countable union of equivalence relations with finite orbits. For our purposes, we only need the fact that if $\mathcal{R}$ is a type I equivalence relation (and hence hyperfinite), then any discrete measured groupoid $\mathcal{G}$ with associated equivalence relation $\mathcal{R}$ can be written as such a semi-direct product.
@@
Recall that a discrete countable group $G$ is called $n$-generated for a positive integer $n$ if $G$ can be generated by $n$ elements as a group. B.H. Neumann constructed an uncountable family of discrete countable groups that are mutually non-isomorphic and are all 2-generated in [Neumann37]. For a reference, we direct the reader to [delaHarpe2000, Section III.B]. We recall the construction here. For all integers $n\geq 1$, let $A_{n}$ denote the alternating group on $n$ elements. Of course the number of elements in $A_{n}$ is $\frac{n!}{2}$ for all $n\geq 1$. It is a standard fact from group theory (for example see [DummitFoote04, Chapter IV.6]) that for $n\geq 5$, the alternating group $A_{n}$ is simple.
@@
Notice that $\alpha_{U}$ is a permutation of infinite order but $\beta_{U}$ has order 3. Let $\mathfrak{G}_{U}$ be the group of permutations of $X_{U}$ generated by $\alpha$ and $\beta$. By analyzing finite normal subgroups of $\mathfrak{G}_{U}$, it turns out that for all $i\geq 1$, $\mathfrak{G}_{U}$ has a finite normal subgroup isomorphic to $A_{u_{i}}$. Moreover, every finite normal subgroup of $\mathfrak{G}_{U}$ is isomorphic to $A_{u_{i}}$ for some $i\geq 1$. For a detailed proof of these facts, we refer the reader to [delaHarpe2000, Points 26 and 27, Chapter III.B]. As an immediate corollary, we have the following proposition.
@@
For two distinct sequences $U$ and $V$ consisting of strictly increasing odd integers $\geq 5$, the 2-generated groups $\mathfrak{G}_{U}$ and $\mathfrak{G}_{V}$ as constructed above are non-isomorphic.

### v2

# Measured groupoids beyond equivalence relations and group actions

@@ 1 Introduction
Discrete measured groupoids were introduced by Mackey in [Mac62] and provide a uniform framework to study discrete countable groups, countable measured equivalence relations and (measurable) actions of countable discrete groups on measure spaces. In [Hah78], Hahn studied the left regular representation of such groupoids and introduced their von Neumann algebras. Measured groupoids are of interest from an operator algebraic point of view because they give a huge class of examples of von Neumann algebras, that covers group von Neumann algebras, von Neumann algebras with Cartan subalgebras and group measure space von Neumann algebras. Every discrete measured groupoid comes equipped with an associated countable measured equivalence relation and a measured field of isotropy groups.
As in the case of groups and group actions, several properties of the associated von Neumann algebra is completely captured by the groupoid. In [BCDK24] we introduced the infinite conjugacy class (icc) property for groupoids and proved that the von Neumann algebra associated to a discrete measured groupoid is a factor if and only if the groupoid is icc and satisfies an ergodicity condition. In some sense, the conditions as well as the proof of the main theorem in [BCDK24] resemble the cases of groups and group actions. Moreover, the literature on discrete measured groupoids lacks concrete examples beyond equivalence relations and transformation groupoids. This naturally leads to the question of how much more general the class of ergodic groupoids is compared to the subclass of ergodic equivalence relations and transformation groupoids arising from ergodic actions. More precisely, are there ergodic discrete measured groupoids that are not isomorphic to equivalence relations or transformation groupoids?
This leads to the natural definition of genuine discrete measured groupoids. We call an ergodic groupoid genuine if it is not isomorphic to a transformation groupoid or an equivalence relation. More generally, a discrete measured groupoid is called genuine if its direct integral decomposition into ergodic groupoids (as in [Hah78]) has a positive measure part consisting of ergodic genuine groupoids. This article deals with the following question.
@@
The main idea comes from the following easy observation: If $G\ltimes X$ is the transformation groupoid of a non-singular group action $G\curvearrowright(X,\mu)$, the isotropy group at a point $x\in X$ is the stabilizer $\{g\in G\;|\;gx=x\}$ of the action, and in particular a subgroup of $G$. Thus if we can construct a measured field of groups $(G_{x})_{x\in X}$ such that they cannot be subgroups of a single discrete countable group, then such a measured field of groups cannot be the isotropy part of a transformation groupoid. To construct such a measured field, we use B.H. Neumann’s construction from [Neu37] of an uncountable family of 2-generated groups that are pairwise non-isomorphic.
The interesting point about Neumann’s construction is that he associates to every strictly increasing sequence of odd integers $U$, a countable set of elements $X_{U}$ and permutations $\alpha_{U},\beta_{U}\in\text{Perm}(X_{U})$ in a way such that the groups $G_{U}$ generated by these permutations are mutually non-isomorphic. We can use the Borel bijection between $(0,1)$ and $\mathbb{N}^{\mathbb{N}}$ to get a standard Borel structure on the set $Y$ of all strictly increasing sequences of odd integers. Due to the canonical nature of the construction, we can endow $(G_{U})_{U\in Y}$ with a standard Borel structure that makes it a Borel field of groups as in Corollary 3.4. Now, by equipping $Y$ with any diffuse Borel probability measure, we get a measured field of groups which can not be the isotropy part of any transformation groupoid.
@@ 2.1 Discrete measured groupoids
The most basic example of a discrete measured groupoid is a countable discrete group with a unit space consisting of a single point. More generally for any such groupoid and any point $x\in X$, the isotropy group at $x$ is the discrete countable group $\Gamma_{x}=\{g\in\mathcal{G}\;|\;s(g)=t(g)=x\}$. The subgroupoid $\operatorname{Iso}(\mathcal{G})$ given by $\{g\in\mathcal{G}\>|\;s(g)=t(g)\}$ is called the isotropy subgroupoid of $\mathcal{G}$ or the isotropy group bundle. The isotropy subgroupoid consists of a measured family of discrete countable groups $(\Gamma_{x})_{x\in X}$, which can be made precise as follows:
@@
    If $\eta$ is a Borel probability measure on $Z$, then $(G_{z})_{z\in Z}$ is called a measured field of discrete countable groups over $(Z,\eta)$ if there is a $\eta$-conull subset $Z_{0}\subseteq Z$ such that with the restricted Borel structure on $\pi^{-1}(Z_{0})$, $(G_{z})_{z\in Z_{0}}$ is a Borel field of discrete countable groups.
We remark here that while defining a Borel fields of Polish groups $K=(K_{x})_{x\in X}$ in the traditional way (see [Sut85]), one also requires the existence of enough Borel sections, i.e., a sequence of Borel sections $\theta_{n}:X\rightarrow K$ such that at each $x\in X$, $\{\theta_{n}({x})\;|\;n\in\mathbb{N}\}$ is dense in $K_{x}$. Such a sequence of sections is called a total set of sections. However note that in the case of discrete countable groups, this condition is automatic, essentially due to the Lusin-Novikov uniformization theorem (see [Kec95, Theorem 18.10]).
@@
Let $(Z,\eta)$ be a standard probability space and let $(G_{z})_{z\in Z}$ be a measured field of discrete countable groups. Let $K$ be a discrete countable group. Then the wreath products $(G_{z}\wr K)_{z\in Z}$ form a measured field of discrete countable groups.
@@
In this case the projections $\pi_{s}$ and $\pi_{t}$ sending an element $(x,y)$ to $y$ and $x$ respectively are countable-to-one maps and send Borel sets to Borel sets. It turns out that the non-singularity condition above is equivalent to $\mu_{s}$ and $\mu_{t}$ being equivalent measures on $\mathcal{R}$ (see [FM77, Theorem 2]). Any discrete measured groupoid $\mathcal{G}$ on $(X,\mu)$ gives rise to a countable measured equivalence relation $\mathcal{R}_{\mathcal{G}}$ given by:
@@
Orbit equivalence relations are the most basic examples of equivalence relations. Let $G$ be a countable discrete group and $G\curvearrowright(X,\mu)$ be a nonsingular action of $G$ on a standard probability space. Then the equivalence relation $\mathcal{R}_{G\curvearrowright X}$ given by $\{(gx,x)\;|\;g\in G,x\in X\}$ is a countable measured equivalence relation called the orbit equivalence relation. As shown in [FM77, Theorem 1], every countable measured equivalence relation is in fact isomorphic to an orbit equivalence relation of a countable group action.
With any such action, one can also associate its transformation groupoid which intersects in some situations with the orbit equivalence relations as we see now. Let $G\curvearrowright(X,\mu)$ be a non-singular action on a standard probability space and consider the groupoid $X\rtimes G=\{(x,g)\;|\;x\in X,g\in G\}$ with unit space $(X,\mu)$ such that $s(x,g)=x$ and $t(x,g)=gx$. Multiplication and inverses are defined naturally as follows for all $g,h\in G$ and $x\in X$:
[…]
The isotropy group $\Gamma_{x}$ at $x\in X$ is then isomorphic to the stabilizer $\{g\in G\;|\;gx=x\}$. If the stabilizer groups are trivial almost everywhere, i.e. when the action is essentially free, then $X\rtimes G$ is isomorphic to the orbit equivalence relation as a discrete measured groupoid, where the isomorphism is given by $(x,g)\mapsto(gx,x)$. However the converse is not true, i.e., there are countable measured equivalence relations which do not arise as orbit equivalence relations of essentially free actions, and are hence not isomorphic to transformation groupoids, as shown by Furman in [Fur99].
@@ 2.2 Equivalence relations and ergodic decomposition
For detailed proofs of most of the facts stated in this section, we refer the reader to [FM77]. Let $\mathcal{R}$ be a countable measured equivalence relation on a standard probability space $(X,\mu)$. Recall that such a countable measured equivalence relation is called ergodic if every $\mathcal{R}$-invariant Borel subset $E\subset X$ is either $\mu$-null or $\mu$-conull.
Recall that a function $f\in L^{\infty}(X,\mu)$ is called $\mathcal{R}$-invariant if $f$ is constant on $\mathcal{R}$-orbits $\mu$-almost everywhere. The space $L^{\infty}(X,\mu)^{\mathcal{R}}$ of $\mathcal{R}$-invariant functions is in fact an abelian von Neumann algebra, which is usually denoted by $L^{\infty}(X,\mu)^{\mathcal{R}}$. It can be checked that $\mathcal{R}$ is ergodic if and only $L^{\infty}(X,\mu)^{\mathcal{R}}=\mathbb{C}\cdot 1$.
Countable measured equivalence relations are broadly classified into three types, analoguous to the classification of von Neumann algebras. Let $\mathcal{R}$ be an ergodic equivalence relation on a standard probability space $(X,\mu)$. Then $\mathcal{R}$ is said to be of type In for $n\in\{1,2,$$\infty\}$ if $\mathcal{R}$ is isomorphic to the complete equivalence relation on a set of $n$ elements. If $\mathcal{R}$ is not necessarily ergodic, then $\mathcal{R}$ is said to be of type In if $\mathcal{R}$ is isomorphic to an equivalence relation on $E\times S$, where $E$ is a standard Borel space and $S$ is a set of $n$ elements given by $(x,k)\sim(y,l)$ if and only $x=y$. An equivalence relation is said to be smooth or discrete or simply of type I if $X$ admits a $\mathcal{R}$-invariant Borel partition $X=\bigsqcup_{n}X_{n}$ such that $\mathcal{R}|_{X_{n}}$ is of type In for $n\in\{1,2,$$\infty\}.$
If $\mathcal{R}$ is not of type $I$, then $\mathcal{R}$ is said to be of type II1 if there is a probability measure $\nu$ on $X$ which is equivalent to $\mu$ and $\mathcal{R}$ is $\nu$-preserving, i.e., $\nu_{s}=\nu_{t}$ on $\mathcal{R}$. If no such equivalent probability measure exists but there is a $\sigma$-finite infinite measure $\nu$ equivalent to $\mu$ such that $\mathcal{R}$ is $\nu$-invariant, then $\mathcal{R}$ is said to be of type II. If there is no invariant finite or infinite measure on $X$ that makes $\mathcal{R}$ invariant, then $\mathcal{R}$ is said to be of type III. If $\mathcal{R}$ is a countable measured equivalence relation on $(X,\mu)$, then $X$ decomposes uniquely (up to null sets) into $\mathcal{R}$-invariant Borel subsets $X_{\text{I}}$, $X_{\text{II}}$ and $X_{\text{III}}$ such that $\mathcal{R}|_{X_{k}}$ is of type $k$ for $k\in\{\text{I, II, III}\}$. For the purposes of this article, we shall only be working with the decomposition of $\mathcal{R}$ into its smooth (type I) and non-smooth (non-type I) part. We fix the following notation:
@@
Type I equivalence relations are precisely the ones which admit a so-called fundamental domain. More precisely we have the following characterization. One can look at standard texts or lecture notes on Borel equivalence relations for a proof, for example see [Cal20, Chapter 2]
@@
By the ergodic decomposition theorem ([Hah78, Theorem 6.1]), every discrete measured groupoid is isomorphic to a direct integral of ergodic discrete measured groupoids. More precisely, given a discrete measured groupoid $\mathcal{G}$, let $\mathcal{G}=\mathcal{G}_{\text{s}}\bigsqcup\mathcal{G}_{\text{ns}}$ be its decomposition into its smooth and non-smooth parts. Then $\mathcal{G}_{\text{ns}}$ is isomorphic to a direct integral of ergodic non-smooth groupoids over some standard Borel space $(Y_{\text{ns}})$ and $\mathcal{G}_{\text{s}}$ admits a decomposition $\mathcal{G}_{\text{s}}=\bigsqcup_{n\in\mathbb{N}}\mathcal{G}_{\text{s}}^{n}$, where $\mathcal{G}_{\text{s}}^{n}$ has associated equivalence relation of type In and $\mathcal{G}_{\text{s}}^{n}$ is a direct integral of ergodic discrete measured groupoids of type In over standard Borel spaces $Y_{n}$. The ergodic decomposition of discrete measured groupoids is furthermore unique up to Borel isomorphisms of the standard Borel spaces $Y_{n}$ and $Y_{\text{ns}}$. For more details we refer the reader to [Hah78, Theorem 6.1]. This motivates the following main definition of this article.
@@ 2.3 Semi-direct product groupoids
Suppose that we have a countable measured equivalence relation $\mathcal{R}$ on a standard probability space $(X,\mu)$ and suppose $G=(G_{x})_{x\in X}$ is a measured field of discrete countable groups. Then an action of $\mathcal{R}$ on $G$ is given by a family of group isomorphisms $\delta_{(y,x)}:G_{x}\rightarrow G_{y}$ for all $(y,x)\in\mathcal{R}$ such that outside a null set we have:
@@
$(h,(z,y))\circ(g,(y,x))$ $=(\delta_{(x,y)}(h)\cdot g,(z,x))$
@@
It was shown in [PSV20, Proposition 6.5] that for every so-called treeable equivalence relation $\mathcal{R}$, any discrete measured groupoid with associated equivalence relation $\mathcal{R}$ can be written as such a semi-direct product. In particular, this is the case if $\mathcal{R}$ is a hyperfinite equivalence relation, i.e., one that can be written as a countable union of equivalence relations with finite orbits. For our purposes, we only need the fact that if $\mathcal{R}$ is a type I equivalence relation (and hence hyperfinite), then any discrete measured groupoid $\mathcal{G}$ with associated equivalence relation $\mathcal{R}$ can be written as such a semi-direct product.
@@
Recall that a discrete countable group $G$ is called $n$-generated for a positive integer $n$ if $G$ can be generated by $n$ elements as a group. B.H. Neumann constructed an uncountable family of discrete countable groups that are mutually non-isomorphic and are all 2-generated in [Neu37]. For a reference, we direct the reader to [Har00, Section III.B]. We recall the construction here. For all integers $n\geq 1$, let $A_{n}$ denote the alternating group on $n$ elements. Of course the number of elements in $A_{n}$ is $\frac{n!}{2}$ for all $n\geq 1$. It is a standard fact from group theory (for example see [DF04, Chapter IV.6]) that for $n\geq 5$, the alternating group $A_{n}$ is simple.
@@
Notice that $\alpha_{U}$ is a permutation of infinite order but $\beta_{U}$ has order 3. Let $\mathfrak{G}_{U}$ be the group of permutations of $X_{U}$ generated by $\alpha$ and $\beta$. By analyzing finite normal subgroups of $\mathfrak{G}_{U}$, it turns out that for all $i\geq 1$, $\mathfrak{G}_{U}$ has a finite normal subgroup isomorphic to $A_{u_{i}}$. Moreover, every finite normal subgroup of $\mathfrak{G}_{U}$ is isomorphic to $A_{u_{i}}$ for some $i\geq 1$. For a detailed proof of these facts, we refer the reader to [Har00, Points 26 and 27, Chapter III.B]. As an immediate corollary, we have the following proposition.
@@
([Har00, Chapter III.B, Proposition 28]) For two distinct sequences $U$ and $V$ consisting of strictly increasing odd integers $\geq 5$, the 2-generated groups $\mathfrak{G}_{U}$ and $\mathfrak{G}_{V}$ as constructed above are non-isomorphic.